# HybridVLA: Collaborative Diffusion and Autoregression in a Unified Vision-Language-Action Model

### v2

@@ 4.1 Simulation Experiment
Ex6 - - 0.10
Ex7 - - - 0.57
Table 3: Impact of each component. AR and Dif represent autoregressive and diffusion-based action generation, respectively, using LLM as backbone. $L_{Hybrid}$ means jointly training the model by Hybrid objectives. LSP denotes large-scale pretraining on assembled robotic datasets, while RSE refers to our injected robot state embedding. CTR and CAE mean our proposed collaborative training recipe and collaborative action ensemble method.
4.2 Ablation Study
We perform ablation experiments on the 10 RLBench tasks and compute the average manipulation accuracy. The impact of each component. As shown in Table 3, a comparison between Ex1 and Ex2, as well as Ex3 and Ex4, demonstrates that under our proposed collaborative training of diffusion and autoregressive action generation, HybridVLA-dif (Ex1) and HybridVLA-ar (Ex2) outperform their individually trained counterparts. These results further confirm that our method effectively integrates diffusion’s continuous prediction capability with the reasoning ability of the autoregressive approach, enabling mutual reinforcement. Compared to Ex0, Ex5 and Ex6 highlight the essential role of large-scale pretraining and robot state injection in ensuring stable prediction. A comparison of Ex7 and Ex1 shows that simply combining diffusion and autoregressive approaches in a single LLM without our training recipe fails to improve accuracy. Moreover, Ex7 underperforms compared to training diffusion alone without hybrid loss (Ex2). The various token formulation designs used in our training recipe are explored in Table 1. The impact of denoising steps. In Figure 3, we explore the relationship between the manipulation performance and different denoising steps, comparing Ex1 (HybridVLA-dif), Ex2, and Ex5. The results suggest that integrating the diffusion process into the LLM and leveraging large-scale robotic pretraining allows for a reduction in inference denoising steps from 20 to 4 without causing performance degradation. Conversely, Ex5 (without large-scale pretraining) exhibits a significant decline in manipulation accuracy. To balance inference speed and accuracy, we set the diffusion denoising steps to 4. In Appendix B.2, we provide additional ablation studies on (1) confidence thresholds in the collaborative action ensemble and (2) KV cache influence on inference speed.
@@ 4.2 Ablation Study
Pick Unplug Pour Wipe Open drawer Mean. Pick Lift ball Place bottles Wipe Fold Mean.
Models and place charger water blackboard and place inside S.R. $\uparrow$ and place and place at rack blackboard shorts S.R. $\uparrow$
$\pi_{0}$ (2.6B) [8] 0.50 0.35 0.45 0.35 0.60 0.45 0.75 0.65 0.40 0.30 0.65 0.55
CogACT (7B) [52] 0.80 0.70 0.40 0.65 0.50 0.61 - - - - - -
@@
Table 4: Comparison of our method and baselines in real-world scenarios. We train all methods in a single-task setting [110] and report the success rates. Success is determined by human evaluation based on whether the task is completed. Since CogAct lacks support for multi-view images, which are crucial for dual-arm tasks [8, 20], we conduct our dual-arm comparison solely with $\pi_{0}$.
4.3 Real-World Experiment
Self-collected Data. For single-arm tasks, we use a Franka Research 3 robot with a static front-view and a wrist-view camera. We perform 5 tasks: 1) Pick and place, 2) Unplug charger, 3) Open drawer and place inside, 4) Pour water, 5) Wipe blackboard. For each task, 100 demonstrations are collected via teleoperation using a SpaceMouse from various positions on the table. For dual-arm tasks, we use an AgileX dual-arm robot equipped with a static exterior view, a right-wrist view, and a left-wrist view camera. We conduct 5 coordinated dual-arm tasks: 1) Pick and place, 2) Lift ball and place, 3) place two bottles at rack, 4) Wipe blackboard, 5) Fold shorts. Similarly, 100 demonstrations are collected for each task using master-puppet teleoperation. Additional details are provided in Appendix A.2.
Training and Evaluation Details. For a fair comparison, we evaluate only the diffusion-generated actions of our method (HybridVLA-dif) against diffusion-based VLA methods, $\pi_{0}$[8] and CogAct[52]. The implementation details remain consistent with our simulation experiments, except for incorporating two-view inputs for single-arm tasks and three-view inputs for dual-arm tasks. For evaluation, we use the checkpoint from the latest epoch to perform 20 rollouts across diverse tabletop positions.
Quantitative Results. In Table 4, HybridVLA-dif achieves outstanding performance across single-arm real-world tasks. For Pick and place and Unplug charger, our method achieves success rates of 85% and 95%, respectively, demonstrating accurate object position prediction. For Pour Water, HybridVLA-dif outperforms the previous SOTA method by 30%, demonstrating its ability to comprehend object relationships and predict precise rotations. The superior performance in Wipe Blackboard and Open Drawer and Place Inside further underscores HybridVLA-dif’s stability in long-horizon tasks. For dual-arm tasks, we extend the action dimension of diffusion tokens to 14-DOF, representing the 7-DOF end-effector poses for both the right and left arms. Our method consistently surpasses previous VLA method across 5 distinct tasks, highlighting HybridVLA’s ability to leverage VLMs’ reasoning capabilities for object affordance understanding and dual-arm coordination.
Qualitative Results. In Figure 4, we present visualizations of the manipulation process performed by our method. HybridVLA-dif accurately predicts control actions across various task demands, including precise positioning and rotation, dual-arm coordination, and scene understanding. For instance, in the Wipe blackboard dual-arm task, our method enables the left arm to precisely grasp the eraser while the right arm holds the edge of the blackboard. It then coordinates the movement and rotation of both objects to establish a relative spatial relationship. Finally, the left arm controls the eraser to precisely remove the red circle. Additional failure case analyses are provided in Appendix D, and execution videos are available in the supplementary materials.
Figure 4: The qualitative results of HybridVLA-dif in real-world tasks. More visualizations are shown in Appendix C.
Task Pick and place(single arm) Lift ball and place(dual arm)
Scenario HybridVLA-dif Cogact HybridVLA-dif $\pi_{0}$
Original 0.85 0.80 0.75 0.65
Object 0.50(-41%) 0.45(-43%) 0.70(-7%) 0.60(-8%)
Background 0.60(-31%) 0.50(-37%) 0.60(-20%) 0.50(-23%)
Height 0.70(-17%) 0.50(-37%) 0.55(-20%) 0.45(-31%)
Lightning 0.70(-17%) 0.60(-25%) 0.65(-13%) 0.55(-15%)
Table 5: Generalization. “Object”, “Background”, “Height”, and “Lighting” denote unseen manipulated objects, backgrounds, spatial positions, and lighting conditions, respectively. The image above illustrates the four unseen test scenarios, with red boxes highlighting the key differences.
@@ 4.4 Generalization Experiment
Beyond achieving stable manipulation across diverse tasks, HybridVLA also exhibits strong generalization capabilities in real-world scenarios. Since CogAct and $\pi_{0}$ excel in single-arm and dual-arm tasks, respectively, we design four common generalization experiments, comparing our method with CogAct on the single-arm Pick and place task and with $\pi_{0}$ on the dual-arm Lift ball and place task. 1) Unseen manipulated objects. In this scenario, we replace the training manipulated objects with a series of unseen objects, e.g., replacing the red block with a charger. As shown in the “Object” row of Table 5, our method demonstrates the smallest accuracy drop. These results indicate that HybridVLA effectively incorporates diffusion into next-token prediction, leveraging VLM pretraining knowledge to reason about diverse object semantics based on observation and language context. 2) Unseen background. In this scenario, cluttered backgrounds are introduced during testing, such as adding unseen flowers around the manipulated object. As shown in the “Background” row of Table 5, HybridVLA exhibits minimal accuracy degradation. These results demonstrate that our collaborative training recipe effectively absorbs the strengths of both action generation methods, enhancing robustness to environmental variations. 3) Unseen Spatial position. Unlike position shifts within the same plane, we introduce height variations during testing, further challenging the model’s spatial comprehension. As shown in the “Height” row of Table 5, HybridVLA consistently achieves precise manipulation even when encountering objects in previously unseen spatial positions. These results highlight that HybridVLA not only excels in observational generalization but also exhibits strong trajectory generalization capabilities. 4) Unseen lighting conditions. Finally, we introduce variations in lighting conditions, a common challenge in real-world environments. As shown in the last row of Table 5, all methods maintain satisfactory performance. This can be attributed to large-scale pretraining on robotic datasets, which significantly enhances these models’ ability to generalize across diverse data distributions.
@@ 5 Conclusion and Limitation
In this paper, we introduce HybridVLA, a unified framework that equips a single LLM with both diffusion and autoregressive action prediction capabilities. Our proposed collaborative training recipe bridges the gap between these two action generation approaches, enabling mutual reinforcement and enhancing manipulation robustness. By effectively integrating the inherently continuous nature of actions from diffusion with the reasoning capabilities of autoregressive methods, HybridVLA achieves outstanding performance and strong generalization across both simulation and real-world tasks. One limitation of HybridVLA is that its inference time is constrained by the slower autoregressive generation, reducing control frequency, similar to previous autoregressive-based VLA methods [47, 10, 56]. However, our collaborative training enables mutual reinforcement between the two generation methods, allowing inference with only the diffusion process (HybridVLA-dif) and achieving a 9.4 Hz inference speed.
@@ A.1 Large-scale Pretraining Dataset
Our pre-training dataset collection comprises 35 datasets, encompassing a total of 760k trajectories and 33m frames. Table 6 provides a comprehensive list of our pre-training datasets along with their respective sampling weights. The number of trajectories and the sampling weights can be automatically adjusted during dataset assembly. Following the prior data preprocessing approach [47], we reformulate the pre-training datasets to emphasize end-effector sequence control, ensuring alignment with the specific requirements of our model training. Due to inherent differences among datasets, only single 2D observations are used during pre-training. However, during fine-tuning, HybridVLA can accommodate both single- and multi-view observations depending on the task requirements. For instance, AgileX dual-arm robot tasks require three viewpoints—an ego view and two wrist camera views—to capture a comprehensive observation of the object while mitigating occlusions caused by the robot arm. HybridVLA processes multi-view images using a shared vision encode and then concatenates the visual feature along the token dimension. Notably, the difference in the number of images used during pre-training and fine-tuning does not impact manipulation performance in downstream tasks.
@@
Fractal [9] 9.1%
Kuka [43] 27.8%
Bridge[19, 95] 4.1%
Taco Play [81, 66] 2.1%
Jaco Play [17] 0.3%
Berkeley Cable Routing [62] 0.2%
Roboturk [65] 1.7%
Viola [120] 0.7%
Berkeley Autolab UR5 [13] 0.9%
Toto [117] 1.5%
Language Table [64] 3.1%
Stanford Hydra Dataset [6] 3.2%
Austin Buds Dataset [119] 0.2%
NYU Franka Play Dataset [16] 0.6%
Furniture Bench Dataset [29] 1.8%
UCSD Kitchen Dataset [108] $<$0.1%
Austin Sailor Dataset [68] 1.6%
Austin Sirius Dataset [57] 1.2%
DLR EDAN Shared Control [76] $<$0.1%
IAMLab CMU Pickup Insert [83] 0.7%
UTAustin Mutex [85] 1.6%
Berkeley Fanuc Manipulation [118] 0.6%
CMU Stretch [67] 0.1%
BC-Z [37] 5.4%
FMB Dataset [63] 5.0%
DobbE [84] 1.0%
DROID [46] 7.2%
Stanford Kuka Dataset [49] 0.1%
Stanford Robocook Dataset [86] 0.1%
Maniskill [26] 6.3%
Berkeley RPT [78] 0.1%
QUT Dexterous Manipulation [11] 0.1%
RoboSet [48] 1.8%
BridgeData V2 [95] 4.7%
RoboMind [101] 5.2%
@@ A.2 Self-collected Real-world Dataset
The experimental assets and environments for the single-arm and dual-arm setups are shown in Figure 5 (a) and (b), respectively. For the single-arm setup, a 3D-printed UMI gripper [15] is attached to the Franka robot and is used across all baselines. We utilize RealSense 435 and RealSense 515 cameras to capture both wrist and front views. For the dual-arm setup, two Orbbec DABAI cameras are used to capture the left and right wrist views, while a RealSense 515 is mounted overhead to capture a static third-person view. We provide a detailed explanation of the real-world tasks and their success conditions. We begin by describing the single-arm tasks:
@@ B.1 Additional Simulation Experiments
In Table 7, we validate the first observed phenomenon mentioned in Section 3.3: different action types within our proposed framework exhibit varying performance across tasks. Meanwhile, we categorize our method into three modes: HybridVLA (7B), HybridVLA-ar (7B), and HybridVLA-dif (7B). All modes undergo joint training using our proposed collaborative training recipe; however, HybridVLA-ar and HybridVLA-dif rely exclusively on autoregressive-based and diffusion-based action generation during inference, respectively. The experiments are conducted in the RLBench simulator across 10 tasks, and evaluated based on success rate. Comparing HybridVLA-ar and HybridVLA-dif, HybridVLA-ar outperforms in 4 out of 10 tasks, while HybridVLA-dif leads in the remaining 6 tasks. These results validate our findings that, within the HybridVLA framework, diffusion-based predictions excel in precise manipulation tasks, such as Phone on base, Toilet seat down, and Close laptop lid, whereas autoregressive predictions perform better in tasks requiring scene-level semantic reasoning, such as Sweep to dustpan, Water plants, and Frame off hanger. Therefore, while collaborative training allows diffusion-based and autoregressive-based action generation to reinforce each other, assembling both methods results in more robust actions.

### v3

@@ 4.2 Ablation Study
Ex6 - 0.68
Table 4: Generalization. “Object", “Background", “Height", and “Lighting" denote unseen manipulated objects, backgrounds, spatial positions, and lighting conditions, respectively. The image above depicts the unseen test scenarios, with red boxes marking the key differences.
Task Pick and place(single arm) Lift ball and place(dual arm)
Scenario HybridVLA Cogact HybridVLA $\pi_{0}$
Original 0.90 0.80 0.80 0.65
Object 0.60(-33%) 0.45(-43%) 0.75(-6%) 0.60(-8%)
Background 0.80(-11%) 0.50(-37%) 0.60(-25%) 0.50(-23%)
Height 0.75(-17%) 0.50(-37%) 0.60(-25%) 0.45(-31%)
Lightning 0.70(-22%) 0.60(-25%) 0.75(-6%) 0.55(-15%)
4.3 Real-World Experiment
Self-collected Data. For single-arm tasks, we use a Franka Research 3 robot with a static front-view and a wrist-view camera. We perform 5 tasks: 1) Pick and place, 2) Unplug charger, 3) Open drawer and place inside, 4) Pour water, 5) Wipe blackboard. For each task, 100 demonstrations are collected via teleoperation using a SpaceMouse device. For dual-arm tasks, we use an AgileX dual-arm robot equipped with a static exterior view, a right-wrist view, and a left-wrist view camera. We conduct 5 coordinated dual-arm tasks: 1) Pick and place, 2) Lift ball and place, 3) place two bottles at rack, 4) Wipe blackboard, 5) Fold shorts. Similarly, 100 demonstrations are collected for each task using master-puppet teleoperation. Additional details are provided in Appendix  A.2.
Training and Evaluation Details. We evaluate HybridVLA (7B) and HybridVLA-dif (7B) against previous VLA methods, $\pi_{0}$ black2024pi_0 and CogAct li2024cogact . The implementation details remain consistent with our simulation experiments, except for using two-view inputs for single-arm tasks and three-view inputs for dual-arm tasks. For evaluation, we use the checkpoint from the latest epoch to perform 20 rollouts across diverse tabletop positions.
Quantitative and Qualitative Results. In Table 5, HybridVLA and HybridVLA-dif achieve outstanding performance across single-arm tasks. For Pick and place and Unplug charger, HybridVLA achieves success rates of 90% and 95%, respectively, demonstrating accurate object position prediction. For Pour water, HybridVLA and HybridVLA-dif outperform the previous SOTA method by 35% and 30%, respectively, showcasing their ability to comprehend object relationships and predict precise rotations. The superior performance on Wipe blackboard and Open drawer and place inside further underscores the robustness of our method in long-horizon tasks. For dual-arm tasks, we extend the action dimensions of both diffusion and autoregressive tokens to 14-DOF, representing the 7-DOF end-effector poses for both the right and left arms. Our method consistently outperforms previous VLA approaches across five distinct tasks, highlighting HybridVLA’s ability to effectively leverage VLMs’ reasoning capabilities for dual-arm coordination in complex scenarios. Furthermore, in the lower part of Table 5, we present visualizations of the manipulation processes performed by our method, which accurately predicts actions across various task demands, including precise positioning and rotation, dual-arm coordination, and scene understanding. Additional qualitative results and failure case analyses are provided in Appendix  C and Appendix  D, respectively, and execution videos are available in the supplementary materials.
Table 5: Real-world experiments. All methods are trained in a single-task setting ze20243d , with success determined by human evaluation. Since CogAct lacks support for multi-view images, which are crucial for dual-arm tasks black2024pi_0 ; fu2024mobile , we conduct our dual-arm comparison solely with $\pi_{0}$.
@@ 4.3 Real-World Experiment
Models Pick Unplug Pour Wipe Open drawer Mean. Pick Lift ball Place bottles Wipe Fold Mean.
and place charger water blackboard and place inside S.R. $\uparrow$ and place and place at rack blackboard shorts S.R. $\uparrow$
$\pi_{0}$ (2.6B) black2024pi_0 0.50 0.35 0.45 0.35 0.60 0.45 0.75 0.65 0.40 0.30 0.65 0.55
CogACT (7B) li2024cogact 0.80 0.70 0.40 0.65 0.50 0.61 - - - - - -
@@
HybridVLA(7B) 0.90 0.95 0.80 0.85 0.65 0.83 0.90 0.80 0.60 0.55 0.70 0.71
@@ 4.4 Generalization Experiment
Since CogAct and $\pi_{0}$ excel in single-arm and dual-arm tasks, respectively, we design four common generalization experiments, comparing our HybridVLA with CogAct on the single-arm Pick and place task and with $\pi_{0}$ on the dual-arm Lift ball and place task. 1) Unseen manipulated objects. In this scenario, we replace the training manipulated objects with a series of unseen objects, e.g., replacing the red block with a charger. As shown in the “Object" row of Table 4.2, our method demonstrates the smallest accuracy drop. These results indicate that, unlike previous diffusion-based VLA methods, HybridVLA effectively integrates diffusion into the autoregressive next-token prediction process, not only capturing the continuous characteristics of diffusion-based generation, but also preserving the object-level semantic reasoning capabilities of autoregressive generation. 2) Unseen background. In this scenario, cluttered backgrounds are introduced during testing, such as adding unseen flowers around the manipulated object. HybridVLA still shows satisfactory results, further demonstrating that our collaborative training recipe effectively inherits the VLM’s scene-level reasoning capabilities, enhancing robustness to environmental variations. 3) Unseen Spatial position. Unlike position shifts within the same plane, we introduce height variations during testing, further challenging the model’s spatial comprehension. As shown in the “Height" row of Table 4.2, HybridVLA consistently achieves precise manipulation even when encountering objects in previously unseen spatial positions. These results highlight that HybridVLA exhibits strong trajectory generalization capabilities through the ensemble of two action generation methods. 4) Unseen lighting conditions. Finally, we introduce variations in lighting conditions, a common challenge in real-world environments. All methods maintain satisfactory performance, demonstrating that large-scale pretraining on robotic datasets enhances their generalization across diverse data distributions.
@@ 5 Conclusion and Limitation
In this paper, we introduce HybridVLA, a unified Vision-Language-Action (VLA) framework that equips a single LLM with both diffusion-based and autoregressive action generation capabilities. To bridge the gap between these two paradigms, we propose a collaborative training recipe that integrates diffusion denoising into the next-token prediction process, enabling mutual reinforcement and improving manipulation robustness. By effectively absorbing the continuous nature of diffusion-based action generation and the semantic reasoning capabilities of autoregressive methods, HybridVLA achieves outstanding performance and strong generalization across both simulation and real-world tasks. One limitation of HybridVLA is that its inference speed is constrained by the slower autoregressive generation, similar to prior autoregressive VLA methods kim2024openvla ; brohan2023rt ; li2024manipllm . However, our collaborative training enables mutual reinforcement between the two generation methods, allowing inference using only the diffusion process (HybridVLA-dif), achieving a 9.4 Hz inference speed. Finally, we state the broader impact of our work in Appendix  E.
@@
Appendix E. A brief conclusion and hope to our work’s broader impact.
@@ A.1 Large-scale Pretraining Dataset
Our pre-training dataset collection comprises 35 datasets, encompassing a total of 760k trajectories and 33m frames. Table 6 provides a comprehensive list of our pre-training datasets along with their respective sampling weights. The number of trajectories and the sampling weights can be automatically adjusted during dataset assembly. Following the prior data preprocessing approach kim2024openvla , we reformulate the pre-training datasets to emphasize end-effector sequence control, ensuring alignment with the specific requirements of our model training. Due to inherent differences among datasets, only single 2D observations are used during pre-training. However, during fine-tuning, HybridVLA can accommodate both single- and multi-view observations depending on the task requirements. For instance, AgileX dual-arm robot tasks require three viewpoints—an ego view and two wrist camera views—to capture a comprehensive observation of the object while mitigating occlusions caused by the robot arm. HybridVLA processes multi-view images using a shared vision encode and then concatenates the visual feature along the token dimension. Notably, the difference in the number of images used during pre-training and fine-tuning does not impact manipulation performance in downstream tasks.
@@
Fractal (brohan2022rt, ) 9.1%
Kuka (kalashnikov2018qt, ) 27.8%
Bridge(ebert2021bridge, ; walke2023bridgedata, ) 4.1%
Taco Play (rosetebeas2022latent, ; mees2023grounding, ) 2.1%
Jaco Play (dass2023jacoplay, ) 0.3%
Berkeley Cable Routing (luo2023multistage, ) 0.2%
Roboturk (DBLP:journals/corr/abs-1811-02790, ) 1.7%
Viola (zhu2023viola, ) 0.7%
Berkeley Autolab UR5 (BerkeleyUR5Website, ) 0.9%
Toto (zhou2023train, ) 1.5%
Language Table (lynch2023interactive, ) 3.1%
Stanford Hydra Dataset (belkhale2023hydra, ) 3.2%
Austin Buds Dataset (zhu2022bottom, ) 0.2%
NYU Franka Play Dataset (cui2022play, ) 0.6%
Furniture Bench Dataset (heo2023furniturebench, ) 1.8%
UCSD Kitchen Dataset (ucsd_kitchens, ) $<$0.1%
Austin Sailor Dataset (nasiriany2022sailor, ) 1.6%
Austin Sirius Dataset (liu2022robot, ) 1.2%
DLR EDAN Shared Control (quere_shared_2020, ) $<$0.1%
IAMLab CMU Pickup Insert (saxena2023multiresolution, ) 0.7%
UTAustin Mutex (shah2023mutex, ) 1.6%
Berkeley Fanuc Manipulation (fanuc_manipulation2023, ) 0.6%
CMU Stretch (mendonca2023structured, ) 0.1%
BC-Z (jang2022bc, ) 5.4%
FMB Dataset (luo2024fmb, ) 5.0%
DobbE (shafiullah2023dobbe, ) 1.0%
DROID (khazatsky2024droid, ) 7.2%
Stanford Kuka Dataset lee2019makingsensevisiontouch 0.1%
Stanford Robocook Dataset shi2023robocooklonghorizonelastoplasticobject 0.1%
Maniskill gu2023maniskill2unifiedbenchmarkgeneralizable 6.3%
Berkeley RPT radosavovic2023real 0.1%
QUT Dexterous Manipulation ceola2023lhmanip 0.1%
RoboSet kumar2023robohive 1.8%
BridgeData V2 walke2023bridgedata 4.7%
RoboMind wu2024robomind 5.2%
@@ A.2 Self-collected Real-world Dataset
The experimental assets and environments for the single-arm and dual-arm setups are shown in Figure 3 (a) and (b), respectively. For the single-arm setup, a 3D-printed UMI gripper chi2024universal is attached to the Franka robot and is used across all baselines. We utilize RealSense 435 and RealSense 515 cameras to capture both wrist and front views. For the dual-arm setup, two Orbbec DABAI cameras are used to capture the left and right wrist views, while a RealSense 515 is mounted overhead to capture a static third-person view. We provide a detailed explanation of the real-world tasks and their success conditions. We begin by describing the single-arm tasks:
@@ B.1 Additional Simulation Experiments
In Table 7, we validate the first observed phenomenon mentioned in Section  3.3: different action types within our proposed framework exhibit varying performance across tasks. Meanwhile, we categorize our method into three modes: HybridVLA (7B), HybridVLA-ar (7B), and HybridVLA-dif (7B). All modes undergo joint training using our proposed collaborative training recipe; however, HybridVLA-ar and HybridVLA-dif rely exclusively on autoregressive-based and diffusion-based action generation during inference, respectively. The experiments are conducted in the RLBench simulator across 10 tasks, and evaluated based on success rate. Comparing HybridVLA-ar and HybridVLA-dif, HybridVLA-ar outperforms in 4 out of 10 tasks, while HybridVLA-dif leads in the remaining 6 tasks. These results validate our findings that, within the HybridVLA framework, diffusion-based predictions excel in precise manipulation tasks, such as Phone on base, Toilet seat down, and Close laptop lid, whereas autoregressive predictions perform better in tasks requiring scene-level semantic reasoning, such as Sweep to dustpan, Water plants, and Frame off hanger. Therefore, while collaborative training allows diffusion-based and autoregressive-based action generation to reinforce each other, assembling both methods results in more robust actions.
HybridVLA(7B) 0.85 0.95 1.0 0.90 1.0 0.50 0.50 0.70 0.50 0.50 0.74
Threshold 0.90 0.92 0.94 0.96 0.98
Success rate 0.66 0.64 0.70 0.74 0.69
B.2 Additional Ablation Study
The impact of confidence threshold in collaborative action ensemble. The proposed collaborative ensemble strategy determines whether to use the action predicted by diffusion alone or the averaged output of both diffusion and autoregressive methods, guided by a mean confidence threshold derived from the autoregressive action token. In this experiment, we investigate the optimal confidence threshold required to ensure the accuracy of autoregressive actions and enhance the overall precision of the ensemble-generated action. Specifically, as shown in Table  8, we vary the threshold from 0.90 to 0.98. We find that when the confidence threshold drops below 0.94, autoregressive predictions become unreliable, leading to a slight degradation in the performance of the ensemble action. Conversely, when the threshold reaches 0.98, the number of valid autoregressive actions becomes too limited, causing the performance of the ensemble action to closely match that of the diffusion-predicted action. Empirically, we conclude that setting the threshold to 0.96 ensures a stable action ensemble.
The impact of KV cache in inference speed. As described in Section  3.3, we adopt the KV cache to eliminate redundant computations and improve inference speed. In this experiment, we examine the extent to which this mechanism accelerates inference. With the KV cache enabled (Table  2 of the main paper), HybridVLA-dif achieves an average success rate of 66% across 10 simulation tasks with an inference speed of 9.4 Hz. Removing it results in a similar average success rate but reduces the inference speed to 5.0 Hz. Although the KV cache has typically been used in previous autoregressive VLA methods kim2024openvla ; li2024manipllm , we are the first to integrate it into an LLM’s diffusion-based action generation.
The impact of denoising steps. In Figure 4, we explore the relationship between manipulation performance and different denoising steps on HybridVLA-dif. Consistent with the findings of previous work bjorck2025gr00t ; liu2024rdt , we reduced the number of DDIM denoising steps of inference from 30 to 4 without observing a significant degradation in manipulation performance. To balance inference speed and accuracy, we set the diffusion denoising steps to 4 in our final implementation.